Question 19 (1 point) Ten years ago, an account was opened with an investment of $5000. Its balance is now $6000. If interest was compounded annually, what interest rate was earned on the account? (Enter your answers as a decimal rounded to 4 decimal places, not a percentage. For example, enter 0.0843 instead of 8.43%) Your Answer: Answer Hide hint for Question 19 Since FV-PV (1+r)An r (FV/PV)A(1/n) -1

Please show work and explain
Answer #1

FV = PV*(1+i)^10

6000 = 5000*(1+i)^10

1.2^(1/10) = 1+i

1.0184 = 1 + i

i = 1.8399%